a2x2+3(3-2ax)= (ax-6)2
Answer by stanbon(75887)

a^2x^2+3(3-2ax)= (ax-6)^2
a^2x^2+9-6ax = a^2x^2-12ax+36

6ax= 27
x=9/2a
Cheers,
Stan H.